Session: Jun 2006 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click

Experiment Feedback

Shanghai (Lingling Wang at Fri Jun 16 16:40:28 2006) Success :

Urumqi (Zhang Hua at Fri Jun 16 16:43:08 2006) Minor failures :
UTC: 8:18 to 12:38 the power is off. No known problems.

Onsala (Michael Lindqvist at Fri Jun 16 16:51:19 2006) Minor failures :
Scan_check failed several times. Note that we had conditioned all the disks before the session. Mark5 log can be found on vlbeer: evn-session2-onsala-part1.mk5.gz rane:/usr2/log/:> grep scan_check ep055on.log | grep ? 2006.166.19:31:02.16/scan_check/46,ep055_180_no0040,?,,,,, 2006.166.23:02:01.50/scan_check/42,ep055_j1746+2603_no0096,mark4,64,2006y166d22h59m59.101s,?,?,? 2006.167.02:39:35.94/scan_check/9,ep055_j1350+3034_no0159,?,,,,, 2006.167.03:54:12.33/scan_check/29,ep055_j2308+0946_no0179,mark4,64,2006y167d3h52m9.150s,?,?,? 2006.167.12:17:11.00/scan_check/6,ep055_j1656+6012_no0313,mark4,64,2006y167d12h15m7.155s,?,?,? 2006.167.15:23:52.72/scan_check/50,ep055_j1327+2210_no0357,mark4,64,2006y167d15h21m49.145s,?,?,? 2006.167.15:58:56.62/scan_check/60,ep055_j1327+2210_no0367,mark4,64,2006y167d15h56m53.164s,?,?,? Disk used in ep055: rane:/usr2/log/:> grep bank_ ep055on.log 2006.166.16:00:39.43&setup01/bank_check 2006.166.16:00:43.98/bank_check/JIVE-067/2000/1024, 2006.166.20:18:28.82/bank_check/JIVE-066/2000/1024, 2006.167.02:09:00.86/bank_check/JIVE-062/2000/1024, 2006.167.08:04:30.32/bank_check/OSOD-007/1317/1024, 2006.167.11:58:17.39/bank_check/MED-0028/3200/1024,

Noto (P. Cassaro at Sat Jun 17 12:34:58 2006) Minor failures : RECO ANT
From 21:25 to 22:07 UT scans lost due to a faulty disk pack. From 15:38 to 15:48 UT antenna stopped for prelimit.

Torun (K. Borkowski at Mon Jun 19 06:23:57 2006) Success :
No known problems.

Westerbork (Tony Foley at Mon Jun 19 08:02:44 2006) Minor failures :
Wb used in phased array mode. scan_check reported lost bytes on scans 0050, 0051,0190,0222,0277,0364,0365. Also gives "?,?,?" for scan 60. Other scan_checks OK

Effelsberg (A. Polatidis at Mon Jun 19 14:44:32 2006) Success :
Severe shadowing by hills

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:27:35 2006) Minor failures : RECO
The MkII telescope was used (Jb2). Mk5 communications lost at 23:15 UT, re-established at 23:56 UT. No other known problems.

Cambridge (Alastair Gunn at Tue Jun 20 09:28:19 2006) Minor failures : RECO
Data throttling reported on disk packs JOD-0021 and JIVE-061, although 1Gbps tests were passed. No replacements were available. No other known problems.

Medicina (Andrea Orlati at Tue Jun 20 09:28:53 2006) Success :
No known problems. Data recording seems to be ok.

EVN experiment feedback facility: entire session

Entire Session

Session: Jun 2006 (session 2)
Wavelengths: 30cm 18/21 5 6 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Jun 5 12:32:29 2006) Minor failures :
This note concerns the UHF experiments only. Shortly before the first experiment we could manage to install the Lange coupler to produce two circular polarizations instead of the linear polarization announced in the block schedule. Also, we failed to install the phase cal.

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:14:12 2006) Success : PCAL
The MkII telescope was used (Jb2). During this session the Mk2 phase cal. signal was found to be too high and could not be easily reduced in strength. I therefore switched off the phas cal. signal for the entire session (for Mk2 only).